A man borrows $5,000 from the bank. Interest is charged at the end of each month 3% per month. He promises to repay at the end of two years the full amount plus interest in one big payment.

i. Calculate the total interest at the end of two years.
ii. How much would be repaid at the end of two years?
iii. If he decides to repay $150 per month, how much will he still owe the bank after the first 2 repayments?
iv. Will he able to repay his debt if he continues to repay at $150 per month?
1
Expert's answer
2018-10-31T11:26:09-0400

i. The total interest at the end of two years is:

TI = 5,000(1 + 0.03)^24 - 5,000 = $5,163.97.

ii. At the end of two years he will repay FV = 5,000(1 + 0.03)^24 = $10,163.97.

iii. If he decides to repay $150 per month, then after the first 2 repayments

he will still owe the bank 10,163.97 - 300 = $9,863.97.

iv. If he continues to repay at $150 per month, then he will pay 150*24 = $3,600, so he won't be able to repay his debt.